OpenStudy (anonymous):

Solve the following equations analytically: 1) 8^((x^3)-x)=1 2) 45((1/5)^x)=10

OpenStudy (anonymous):

first one gives \(x^3-x=0\) or \(x(x^2-1)=0\) making \(x(x+1)(x-1)=0\)

OpenStudy (anonymous):

so \(x=0\) or \(x=1\) or \(x=-1\)

OpenStudy (anonymous):

so what happened to the 8 and the 1 in the first question

hero (hero):

^See, if you try to isolate x, you just end up eliminating it in the process.

OpenStudy (anonymous):

So what should the final answer be?

hero (hero):

Actually, take logs of both sides and see if that helps

hero (hero):

Yes, taking logs of both sides should work

hero (hero):

Because after taking logs of both sides, you'll end up with \[(x^3 - x)\log(8) = \log(1)\]

hero (hero):

Then you'll have: \[(x^3 - x) = \frac{\log(1)}{\log(8)}\]

hero (hero):

But log(1) = 0, so \[x^3 - x = \frac{0}{\log(8)}\] Thus: \[x^3 - x = 0\]

hero (hero):

You should be able to solve from there

OpenStudy (anonymous):

ok i got -1, 0, and 1 for x

hero (hero):

Because \(x^3 - x\) reduces to \(x(x+1)(x-1)\)

hero (hero):

Yes, you have it

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Do you know anything about #2

hero (hero):

Give me a minute

OpenStudy (anonymous):

ok

hero (hero):

1. Divide both sides by 45 2. Reduce 10/45 3. Take logs of both sides

OpenStudy (anonymous):

well, i guess i got it :) thanks for your help